# Inducing Dyslexia in Vision Language Models

### v2

@@ Abstract
Dyslexia, a neurodevelopmental disorder characterized by persistent reading difficulties, is often linked to reduced activity of the visual word form area in the ventral occipito-temporal cortex. Traditional approaches to studying dyslexia, such as behavioral and neuroimaging methods, have provided valuable insights but remain limited in their ability to test causal hypotheses about the underlying mechanisms of reading impairments. In this study, we use large-scale vision-language models (VLMs) to simulate dyslexia by functionally identifying and perturbing artificial analogues of word processing. Using stimuli from cognitive neuroscience, we identify visual-word-form-selective units within VLMs and demonstrate that targeted ablation of these units, unlike ablation of random units, leads to selective impairments in reading tasks while general visual and language comprehension abilities remain intact. In particular, the resulting model matches dyslexic humans’ phonological deficits without a significant change in orthographic processing. Taken together, our modeling results replicate key characteristics of dyslexia and establish a computational framework for investigating reading disorders.  [^1] Code available via GitHub.
@@ 2 Background & Related Work
Dyslexia is characterized by persistent difficulties in accurate and/or fluent word recognition and spelling, despite adequate education, normal intelligence, and intact sensory abilities (Lyon et al., 2003). These difficulties are closely linked to deficits in phonological awareness and processing, often manifesting as problems with decoding unfamiliar or nonsense words, rapid naming, and word finding. Further challenges may include variable difficulties with letter–sound learning, oral reading accuracy, written composition, and reading comprehension (Roitsch & Watson, 2019). Importantly, studies indicate that the core mechanisms of dyslexia are consistent regardless of IQ (Stanovich, 2005; Tanaka et al., 2011).
Hypothesized neural substrate.
@@
A growing body of evidence highlights the central role of the Visual Word Form Area (VWFA) in the neurobiology of dyslexia. The VWFA, located in the left ventral occipitotemporal cortex, is critical for fluent reading as it specializes in the rapid and automatic recognition of written words (Dehaene & Cohen, 2011; McCandliss et al., 2003). In individuals with dyslexia, this region often shows functional and anatomical abnormalities. Longitudinal studies have revealed that dyslexic readers consistently exhibit smaller and less selective VWFAs compared to typical readers, even after significant improvements in reading performance through targeted training (Mitchell et al., 2025; Brem et al., 2020). These persistent differences suggest that VWFA atypicalities are not merely a consequence of poor reading experience but represent a stable neurobiological trait of dyslexia. Empirical studies consistently report VWFA hypoactivation and structural differences in dyslexics across languages and cultures (Paulesu et al., 2001; Silani et al., 2005), reinforcing its cross-linguistic relevance. Moreover, lesions in this region have been causally linked to acquired reading disorders (Turkeltaub et al., 2013), underscoring the VWFA’s essential role in both typical and atypical reading processes. Nevertheless, the causal status of VWFA abnormalities remains debated, with some studies suggesting that these differences may reflect a consequence rather than a primary cause of dyslexia (Olulade et al., 2013; Valdois, 2010).
Models of Brain Function.
@@
Computational neural networks have proven effective at predicting both behavioral performance and neural responses in healthy subjects, making them valuable tools for studying brain function. In the visual domain, deep convolutional as well as transformer-based models have successfully captured neural activity in the primate ventral visual stream across object and scene recognition tasks (e.g., Yamins et al., 2014; Khaligh-Razavi & Kriegeskorte, 2014; Schrimpf et al., 2018; 2020; Cadena et al., 2019; Spoerer et al., 2020; Zhuang et al., 2021; Wang et al., 2023; Margalit et al., 2024; Gokce & Schrimpf, 2024; Lonnqvist et al., 2025; Tang et al., 2025). In the language domain, transformer-based and recurrent models have accurately predicted neural responses related to semantic, syntactic, and phonological processing (e.g., Schrimpf et al., 2021; Caucheteux et al., 2022; Goldstein et al., 2022; Toneva et al., 2018; Hosseini et al., 2024; Aw et al., 2023; Tuckute et al., 2024; Rathi et al., 2024; AlKhamissi et al., 2025).
@@ 3 Benchmarks
is a browser-based, self-administered lexical decision task that evaluates core reading skills by measuring both accuracy and speed in distinguishing real words (e.g., “able”, “animal”) from pseudo words (e.g., “ablood”, “accastanct”) (Yeatman et al., 2021; more examples in Appendix Table LABEL:table:examples). In our model, however, response time is not a consideration, so we evaluate performance solely based on the accuracy of lexical decisions. On each trial, the model is shown an image of a letter string and must decide whether it is a real word or a pseudo word. We present 200 real words and 200 pseudo words drawn from the full ROAR-Word corpus as the train set for finding the minimal VWFA mask, with the remaining 50 real words and 50 pseudo words serving as the test set for lexical evaluation. Dyslexia Threshold: We define 65% ROAR performance as the threshold at which subjects are considered as dyslexic. This number is one standard deviation below the mean ROAR-score of the human population, consistent with prior works in humans (Wagner et al., 2020). This choice is also meaningful from an epidemiological perspective as dyslexia is estimated to affect about 5–20% of people (Wagner et al., 2020), and this prevalence aligns well with scores in the range of $[1.65,0.95]$ standard deviations below the mean.
Raven’s Progressive Matrices (RAVEN)
@@
is a nonverbal assessment of fluid intelligence in which participants must select, from five candidate panels, the one that correctly completes a 3×3 or 2×2 matrix by preserving an underlying pattern of shapes or spatial relations (Burke, 1958; Zhang et al., 2019; examples in Appendix Fig. 7). Because it minimizes linguistic and cultural biases, RAVEN serves as an ideal control for general visual‐spatial reasoning in our study. In our experiments, we administered the easy section of the RAVEN clinical version which consists of twelve items. By assessing the model’s accuracy on these twelve puzzles, we establish a baseline for general visual intelligence against which any reading‐specific impairments can be contrasted.
Kempler’s Sentence Comprehension Test
@@
is a clinical assessment designed to evaluate syntactic comprehension by presenting participants with pairs of images accompanied by a spoken or written sentence (Kempler et al., 1998). The task is to choose the image that correctly corresponds to the sentence, making it a useful probe of grammatical understanding and sentence-level semantics. In our implementation, we adapt this test for model-based evaluation by using only the image pairs from the original stimulus set, while providing the sentence as a text prompt. The model is asked to decide which image matches a given caption (examples in Appendix Fig. 8). This transforms the task into a visual question answering (VQA) challenge, allowing us to assess the model’s capacity for sentence comprehension grounded in visual reasoning.
Lexical Decision with Orthographic and Phonological Manipulations.
@@
To further dissect reading impairments into orthographic and phonological components, we incorporated stimuli from a lexical decision task developed by Luke et al. (2023). This benchmark includes four stimulus types: (1) homophones (e.g., “brake” and “break”), (2) pseudo-homophones (e.g., “beaf” and “birf”), (3) transposed-letter (TL) neighbour words (e.g., “blots” and “bolts”), and (4) TL non-words (e.g., “golve” and “glove”). These stimuli were originally used to contrast phonological and orthographic theories of dyslexia in human readers. For analysis, we group the stimuli into two categories: phonology-sensitive (homophones and pseudo-homophones) and orthography-sensitive (TL words and TL non-words). The model performs a lexical decision on each item (an image containing a word or a non-word), and accuracy serves as the primary outcome measure. By comparing performance across these two groups, we can determine whether the model finds phonologically demanding items or orthographically demanding items more challenging.
@@ 4 Methodology
Functional Localization of VWF-selective Units.
Given the well-established role of the VWFA in dyslexia, we target visual-word-form-selective (VWF-selective) units in vision-language models (VLMs) as regions of interest for artificial lesions. To reflect the dual involvement of visual and linguistic processes in dyslexia, we focus on computational models that process images (and text) as input which reflect the full processing chain from retina (pixel) input to language. These VLMs, by presenting visual stimuli as image input and task instructions as text tokens, allow us to dissociate visual processing impairments from general language deficiencies, and their integration of visual and linguistic modalities within a single architecture provides a unified framework for examining the cortical processes underlying word recognition and reading impairments in dyslexia. Our method enables a hypothesis-driven investigation of reading-specific mechanisms directly within artificial systems and provides a scalable and manipulable alternative to costly human studies.
@@
To induce a dyslexia‐like impairment, we gradually increase the fraction of the top k% VWF–selective units that are ablated (i.e., set to zero) within the model’s language layers. We begin with no masking (0%) and repeatedly mask a larger proportion of these units, evaluating at each level the model’s accuracy on the train subset of ROAR stimuli. We stop increasing the mask size as soon as the ROAR score falls below 65%, which we defined as the ROAR dyslexia threshold (Section 3). The smallest proportion of masked units that meets this criterion defines our VWF–selective units (see also Fig. 3a).
Model Details.
@@
We focused our analyses on Qwen2-VL-72B (Yang et al., 2024), selected for its strong performance across benchmarks, particularly in OCR and visual understanding tasks. Qwen2-VL-72B is a large-scale autoregressive transformer integrating visual and textual modalities. Within the model, we examined the MLP layers of the language decoder where the minimal subnetwork for dyslexia corresponds to approximately 6.89% of all units (see below for hyperparameter analyses). Our approach successfully elicited reading deficits in all tested models (additionally LLaVA-NeXT-13B, Liu et al., 2024; and PixTral-12B, Agrawal et al., 2024). These deficits selectively affected reading only, without significant differences on visual IQ benchmarks as long as baseline performance was significantly above chance (Appendix Fig. 6).
Perturbation Strength.
@@
Beyond a full ablation (activity = 0), we additionally experimented with parametric perturbations of VWFA units by varying the scaling factor on the activations of the selected VWF-selective units. Specifically, we tested scaling factors in the range [2,4][-2,4], including sublinear, neutral (1.0), and superlinear amplification. We observed that only complete ablation (scaling factor = 0) reliably produced the selective impairments characteristic of dyslexia (Fig. 3b). Perturbations at other scales either had negligible behavioral effects or resulted in severely degraded output coherence, producing nonsensical or empty responses (Table 2). Consequently, we focus our analysis on full ablation as the most faithful computational analogue of focal VWFA disruption.
Layer Choice.
@@
To identify VWF-selective units, we evaluated subnetworks from different layer types across the model. The MLP gate projection layers in the language decoder emerged as the most selective, aligning with prior findings that MLP layers exhibit strong knowledge-specific selectivity (Meng et al., 2022; Zhang et al., 2021). To support this interpretation, we ablated minimal subnetworks from other components of the model, the vision encoder (visual.blocks.{i}.attn.proj), the visual merger (visual.merger.mlp.{i}), and the language decoder’s self-attention outputs (model.layers.{i}.self_attn.o_proj). These ablations led to sharper declines in RAVEN performance relative to ROAR performance, indicating that the observed effects are not purely reading-selective (Fig. 3c). Based on these results, we define VWF-selective units as those located in the MLP layers (model.layers.{i}.mlp.gate_proj) across all 80 transformer blocks of the language decoder (Fig. 3d).

### v3

@@ Abstract
Dyslexia, a neurodevelopmental disorder characterized by persistent reading difficulties, is often linked to reduced activity of the visual word form area (VWFA) in the ventral occipito-temporal cortex. Traditional approaches to studying dyslexia, such as behavioral and neuroimaging methods, have provided valuable insights but remain limited in their ability to test causal hypotheses about the underlying mechanisms of reading impairments. In this study, we use large-scale vision-language models (VLMs) to simulate dyslexia by functionally identifying and perturbing artificial analogues of word processing. Using stimuli from cognitive neuroscience, we identify visual-word-form-selective units within VLMs and demonstrate that they predict human VWFA neural responses. Ablating model VWF units leads to selective impairments in reading tasks while general visual and language comprehension abilities remain intact. In particular, the resulting model matches dyslexic humans’ phonological deficits without a significant change in orthographic processing, and mirrors dyslexic behavior in font sensitivity. Taken together, our modeling results replicate key characteristics of dyslexia and establish a computational framework for investigating brain disorders.  [^1] Code available via GitHub.
@@ 2 Background & Related Work
Behavioral effects of dyslexia. Dyslexia is characterized by persistent difficulties in accurate and/or fluent word recognition and spelling, despite adequate education, normal intelligence, and intact sensory abilities (Lyon et al., 2003). These difficulties are closely linked to deficits in phonological awareness and processing, often manifesting as problems with decoding unfamiliar or nonsense words, rapid naming, and word finding. Further challenges may include variable difficulties with letter–sound learning, oral reading accuracy, written composition, and reading comprehension (Roitsch & Watson, 2019). Importantly, studies indicate that the core mechanisms of dyslexia are consistent regardless of IQ (Stanovich, 2005; Tanaka et al., 2011).
@@
Hypothesized neural substrate. A growing body of evidence highlights the central role of the Visual Word Form Area (VWFA) in the neurobiology of dyslexia. The VWFA, located in the left ventral occipitotemporal cortex, is critical for fluent reading as it specializes in the rapid and automatic recognition of written words (Dehaene & Cohen, 2011; McCandliss et al., 2003). In individuals with dyslexia, this region often shows functional and anatomical abnormalities. Longitudinal studies have revealed that dyslexic readers consistently exhibit smaller and less selective VWFAs compared to typical readers, even after significant improvements in reading performance through targeted training (Mitchell et al., 2025; Brem et al., 2020). These persistent differences suggest that VWFA atypicalities are not merely a consequence of poor reading experience but represent a stable neurobiological trait of dyslexia. Empirical studies consistently report VWFA hypoactivation and structural differences in dyslexics across languages and cultures (Paulesu et al., 2001; Silani et al., 2005), reinforcing its cross-linguistic relevance. Moreover, lesions in this region have been causally linked to acquired reading disorders (Turkeltaub et al., 2013), underscoring the VWFA’s essential role in both typical and atypical reading processes. Nevertheless, the causal status of VWFA abnormalities remains debated, with some studies suggesting that these differences may reflect a consequence rather than a primary cause of dyslexia (Olulade et al., 2013; Valdois, 2010).
@@
Models of Brain Function. Computational neural networks have proven effective at predicting both behavioral performance and neural responses in healthy subjects, making them valuable tools for studying brain function. In the visual domain, deep convolutional as well as transformer-based models have successfully captured neural activity in the primate ventral visual stream across object and scene recognition tasks (e.g., Yamins et al., 2014; Khaligh-Razavi & Kriegeskorte, 2014; Schrimpf et al., 2018; 2020; Cadena et al., 2019; Spoerer et al., 2020; Zhuang et al., 2021; Wang et al., 2023; Margalit et al., 2024; Gokce & Schrimpf, 2024; Lonnqvist et al., 2025; Tang et al., 2025). In the language domain, transformer-based and recurrent models have accurately predicted neural responses related to semantic, syntactic, and phonological processing (e.g., Schrimpf et al., 2021; Caucheteux et al., 2022; Goldstein et al., 2022; Toneva et al., 2018; Hosseini et al., 2024; Aw et al., 2024; Tuckute et al., 2024; Rathi et al., 2025; AlKhamissi et al., 2025; Du et al., 2025).
@@ 3 Benchmarks
Rapid Online Assessment of Reading (ROAR) is a browser-based, self-administered lexical decision task that evaluates core reading skills by measuring both accuracy and speed in distinguishing real words (e.g., “able”, “animal”) from pseudo words (e.g., “ablood”, “accastanct”) (Yeatman et al., 2021; more examples in Table LABEL:table:examples). In our model, however, response time is not a consideration, so we evaluate performance solely based on the accuracy of lexical decisions. On each trial, the model is shown an image of a letter string and must decide whether it is a real word or a pseudo word. We present 200 real words and 200 pseudo words drawn from the full ROAR-Word corpus as the train set for finding the minimal VWFA mask, with the remaining 50 real words and 50 pseudo words serving as the test set for lexical evaluation. Dyslexia Threshold: We define 65% ROAR performance as the threshold at which subjects are considered as dyslexic. This number is one standard deviation below the mean ROAR-score of the human population, consistent with prior works in humans (Wagner et al., 2020). This choice is also meaningful from an epidemiological perspective as dyslexia is estimated to affect about 5–20% of people (Wagner et al., 2020), and this prevalence aligns well with scores in the range of $[1.65,0.95]$ standard deviations below the mean.
@@
Raven’s Progressive Matrices (RAVEN) is a nonverbal assessment of fluid intelligence in which participants must select, from five candidate panels, the one that correctly completes a 3×3 or 2×2 matrix by preserving an underlying pattern of shapes or spatial relations (Burke, 1958; Zhang et al., 2019; examples in Appendix Fig. 9). Because it minimizes linguistic and cultural biases, RAVEN serves as an ideal control for general visual‐spatial reasoning in our study. In our experiments, we administered the easy section of the RAVEN clinical version which consists of twelve items. By assessing the model’s accuracy on these twelve puzzles, we establish a baseline for general visual intelligence against which any reading‐specific impairments can be contrasted.
@@
Kempler’s Sentence Comprehension Test is a clinical assessment designed to evaluate syntactic comprehension by presenting participants with pairs of images accompanied by a spoken or written sentence (Kempler et al., 1998). The task is to choose the image that correctly corresponds to the sentence, making it a useful probe of grammatical understanding and sentence-level semantics. In our implementation, we adapt this test for model-based evaluation by using only the image pairs from the original stimulus set, while providing the sentence as a text prompt. The model is asked to decide which image matches a given caption (examples in Appendix Fig. 10). This transforms the task into a visual question answering (VQA) challenge, allowing us to assess the model’s capacity for sentence comprehension grounded in visual reasoning.
@@
Lexical Decision with Orthographic and Phonological Manipulations. To further dissect reading impairments into orthographic and phonological components, we incorporated stimuli from a lexical decision task developed by Luke et al. (2023). This benchmark includes four stimulus types: (1) homophones (e.g., “brake” and “break”), (2) pseudo-homophones (e.g., “beaf” and “birf”), (3) transposed-letter (TL) neighbour words (e.g., “blots” and “bolts”), and (4) TL non-words (e.g., “golve” and “glove”). These stimuli were originally used to contrast phonological and orthographic theories of dyslexia in human readers. For analysis, we group the stimuli into two categories: phonology-sensitive (homophones and pseudo-homophones) and orthography-sensitive (TL words and TL non-words). The model performs a lexical decision on each item (an image containing a word or a non-word), and accuracy serves as the primary outcome measure. By comparing performance across these two groups, we can determine whether the model finds phonologically demanding items or orthographically demanding items more challenging.
Figure 5: VWF-selective units align with human VWFA activity. Alignment between predicted and actual fMRI responses in the human VWFA (noise-normalized), computed across the top-k% of VWF-selective units of the model (dark blue bars) and randomly selected units (light blue bars) for 5 seeds. At small subset sizes, VWF-selective units show significantly higher alignment than the average of five randomly selected subsets. Error bars indicate 95% confidence intervals across seeds.
@@ 4 Methodology
Functional Localization of VWF-selective Units. Given the well-established role of the VWFA in dyslexia, we target visual-word-form-selective (VWF-selective) units in vision-language models (VLMs) as regions of interest for artificial lesions. To reflect the dual involvement of visual and linguistic processes in dyslexia, we focus on computational models that process images (and text) as input which reflect the full processing chain from retina (pixel) input to language. These VLMs, by presenting visual stimuli as image input and task instructions as text tokens, allow us to dissociate visual processing impairments from general language deficiencies, and their integration of visual and linguistic modalities within a single architecture provides a unified framework for examining the cortical processes underlying word recognition and reading impairments in dyslexia. Our method enables a hypothesis-driven investigation of reading-specific mechanisms directly within artificial systems and provides a scalable and manipulable alternative to costly human studies.
@@
Minimal Subnetwork for Dyslexia Simulation. To induce a dyslexia‐like impairment, we gradually increase the fraction of the top k% VWF–selective units that are ablated (i.e., set to zero) within the model’s language layers. We begin with no masking (0%) and repeatedly mask a larger proportion of these units, evaluating at each level the model’s accuracy on the train subset of ROAR stimuli. We stop increasing the mask size as soon as the ROAR score falls below 65%, which we defined as the ROAR dyslexia threshold (Section 3). The smallest proportion of masked units that meets this criterion defines our VWF–selective units (see also Fig. 3a).
@@
Model Details. We focused our analyses on Qwen2-VL-72B (Yang et al., 2024), chosen for its strong OCR and visual understanding performance. Qwen2-VL-72B is a large-scale autoregressive transformer integrating visual and textual modalities. Within the model, we examined the MLP layers of the language decoder where the minimal subnetwork for dyslexia corresponds to approximately 6.89% of all units (see below for hyperparameter analyses). We successfully elicited reading deficits in all tested models (additionally Molmo-72B, Deitke et al., 2024; and PixTral-12B, Agrawal et al., 2024). Deficits were reading-specific, with no significant effects on visual IQ benchmarks (Appendix Fig. 8).
@@
Perturbation Strength. Beyond a full ablation (activity = 0), we additionally experimented with parametric perturbations of VWFA units by varying the scaling factor on the activations of the selected VWF-selective units. Specifically, we tested scaling factors in the range [2,4][-2,4], including sublinear, neutral (1.0), and superlinear amplification. We observed that only complete ablation (scaling factor = 0) reliably produced the selective impairments characteristic of dyslexia (Fig. 3b). Perturbations at other scales either had negligible behavioral effects or resulted in severely degraded output coherence, producing nonsensical or empty responses (Table 3). Consequently, we focus our analysis on full ablation as the most faithful computational analogue of focal VWFA disruption.
@@
Layer Choice. To identify VWF-selective units, we evaluated subnetworks from different layer types across the model. The MLP gate projection layers in the language decoder emerged as the most selective, aligning with prior findings that MLP layers exhibit strong knowledge-specific selectivity (Meng et al., 2022; Zhang et al., 2022). To support this interpretation, we ablated minimal subnetworks from other components of the model, the vision encoder (visual.blocks.{i}.attn.proj), the visual merger (visual.merger.mlp.{i}), and the language decoder’s self-attention outputs (model.layers.{i}.self_attn.o_proj). These ablations led to sharper declines in RAVEN performance relative to ROAR performance, indicating that the observed effects are not purely reading-selective (Fig. 3c). Based on these results, we define VWF-selective units as those located in the MLP layers (model.layers.{i}.mlp.gate_proj) across all 80 transformer blocks of the language decoder (Fig. 3d).
5 Results
Figure 7: Model predicts font-specific reading performance. Accuracy on ROAR lexical decision items across fonts for intact (light blue) and ablated “dyslexic” (dark blue) model, averaged over 10 random seeds. The dyslexic model shows significantly improved performance for OpenDyslexic, OpenDyslexic on an orange background, Comic Sans, and KG Primary Penmanship ($p\ll 0.001$), as well as for Orange Arial ($p\ll 0.01$), while performance is significantly worse for Papyrus ($p\ll 0.01$). In contrast, the intact model’s accuracy remains stable across fonts with no significant differences. Error bars denote 95% confidence intervals across seeds.
Selective Reading Deficits, Preserved Reasoning Capabilities. Ablating visual-word-form (VWF)-selective units leads to a substantial decline in lexical decision accuracy on the ROAR lexical reading task (Fig. 4a). Table LABEL:table:examples shows example responses of the ablated model. Performance in fact dropped below the dyslexia threshold on the held-out test set, a marked reduction from the model’s pre-ablation accuracy ($32\%$ in normalized accuracy with $p\ll 0.01$ using a one-tailed Student’s t-test for whether ablated model performance was significantly lower than baseline performance on each of the tasks; see Appendix A.4 for more details). In contrast, performance on the RAVEN’s Progressive Matrices task remained stable ($p\approx 0.75$), indicating that the lesioned subnetwork selectively impairs reading ability while sparing broader visual reasoning processes. We further evaluated sentence comprehension and visual grounding using the Kempler Sentence Comprehension Test as another visual Q&A task. Following ablation, the model’s accuracy on the Kempler task did not decrease compared to the non-ablated baseline ($p\approx 1$, one-tailed test; a two-tailed Student’s t-test reveals a statistically significant increase in performance of $10\%$, $p\ll 0.01$). This pattern parallels empirical findings in some dyslexic individuals, where deficits in phonological or orthographic processing are accompanied by preserved or even enhanced performance in visual-spatial reasoning and comprehension tasks (von Károlyi et al., 2003; Stanovich, 2005; Tanaka et al., 2011).
Ablating Random Units Does Not Lead to Selective Reading Impairments. To verify that the observed dyslexia-like effects were not simply due to the number or layer-wise distribution of ablated units, but rather depended on the specific identity of the VWF-selective units, we conducted a control experiment. We preserved both the total number of ablated units and the selected layers, but randomly selected which specific units were ablated within each layer. This approach did not lead to selective reading impairments: the performance on all benchmarks dropped ($21\%$ with $p<0.003$ for ROAR, $21\%$ with $p<0.004$ for RAVEN, and $13\%$ with $p<0.042$ for Kempler), and reading performance did not drop below the dyslexia threshold ($p\approx 0.87$; Fig. 4b). Furthermore, when the same number of units was ablated but distributed randomly across the entire network rather than within the same layers, the model’s outputs largely degraded, often generating incoherent or empty responses, confirming that the reading-specific impairment depends on targeting VWF-selective units.